OpenStudy (anonymous):

Use the addition of sines identity to evaluate sin255 degrees what would theta 1 and theta 2 be?

OpenStudy (jdoe0001):

keep in mind that 90 +45 = 135 60+60 = 120 and those are common known angles thus keep in mind that 135 + 120 = 255

OpenStudy (anonymous):

so would I need to solve for 2 different equations?

OpenStudy (anonymous):

one being sin(90+45) and the other sin(60+60) ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... nope notice the instruction "Use the addition of sines identity" thus \(\bf sin({\color{brown}{ \alpha}} + {\color{blue}{ \beta}})=sin({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}}) + cos({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}}) \\ \quad \\ thus \\ \quad \\ sin(255^o)\\ \quad \\ sin({\color{brown}{ 135^o}} + {\color{blue}{ 120^o}})=sin({\color{brown}{ 135^o}})cos({\color{blue}{ 120^o}}) + cos({\color{brown}{ 135^o}})sin({\color{blue}{ 120^o}})\)

OpenStudy (jdoe0001):

and you can get those from your Unit Circle

OpenStudy (anonymous):

Is the answer -1/2? @jdoe0001